Using the column of data HEADLEN, the head length in inches:

  1. Construct a frequency distribution by hand with 5 or 6 classes. Sort the data. Identify the minimum and maximum data and the class information: class width, class limits, class boundaries, and class midpoints. Show all work and the resulting frequency distribution table.   
  2. Hand draw the histogram for the frequency distribution. Make it look professional; title and label appropriately. Use a ruler.   
  3. Describe the shape of the histogram, noting any outliers, gaps, or skewness. What changes might be made to improve the representation of the data?
  4. Identify the four measures of center (using formulas or definitions) and estimate the standard deviation (using the range rule of thumb). You must show your work. Describe the effects of the measures of center and the standard deviation on the shape of the distribution.   
  5. Let Excel choose the classes and do a frequency distribution and histogram for HEADLEN. Compare the results to the frequency distribution and histo gram that you did.   
  6. Let Excel find the descriptive statistics for the column HEADLEN. Compare to the answers you found in question 4.   
  7. An anthropologist found what appears to be a prehistoric bear skull. How could he use Data Set 9 to assess his find?   
  8. Identify at least two other professional fields that might use similar sets of measurements in conducting research
    HEADLEN
    11.0
    16.5
    15.5
    17.0
    15.5
    13.0
    15.0
    13.5
    13.5
    12.5
    16.0
    9.0
    12.5
    14.0
    11.5
    13.0
    13.5
    9.0
    13.0
    16.0
    12.5
    13.0
    13.0
    10.0
    16.0
    10.0
    13.5
    13.0
    13.0
    14.5
    9.5
    13.5
    14.5
    11.0
    11.5
    13.0
    16.5
    14.0
    13.5
    15.5
    11.5
    12.0
    15.5
    9.0
    14.5
    13.5
    10.0
    10.0
    11.5
    11.5
    11.0
    10.0
    15.5
    12.5
Answer #1

minimum = 9
maximum = 17

range = 17-9 = 8

we divide in 5 class
8/5 = 1.6
we can class width = 2
class limits 9 - 10.9 ,11-12.9,..

distribution looks symmetric , there does not seem to any outlier

Please give me a thumbs-up if this helps you out. Thank you! :)